OpenStudy (anonymous):

MEDAL! Jason considered two similar televisions at a local electronics store. The generic version was based off the brand name and was three eighths the size of the brand name. If the generic television set is 12 inches by 24 inches, what are the dimensions of the brand name television? 7.5 inches by 15 inches 28 inches by 56 inches 32 inches by 64 inches 4.5 inches by 9 inches

OpenStudy (anonymous):

@jigglypuff314 some help please

OpenStudy (anonymous):

u their?

OpenStudy (anonymous):

@Mr.ClayLordMath some help

jigglypuff314 (jigglypuff314):

12 times (3/8) = ? 24 times (3/8) = ?

OpenStudy (anonymous):

4.5 9

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

Forget that

OpenStudy (anonymous):

It's simple, the answer is 4.5 inches by 9 inches o.0

OpenStudy (anonymous):

No I was looking at a other question and I got confused

OpenStudy (anonymous):

Ah, i see.

OpenStudy (anonymous):

can u guys help me with one more

OpenStudy (anonymous):

I'll see if i can

OpenStudy (anonymous):

Hexagon IJKLMN is shown on the coordinate plane below.

OpenStudy (anonymous):

If hexagon IJKLMN is dilated by a scale factor of two fifths from the origin to create hexagon I'J'K'L'M'N', what is the ordered pair of point I'? (-0.8, 2.4) (-5, 15) (-2.5, 5) (-0.4, 0.8)

jigglypuff314 (jigglypuff314):

multiply each coordinate point by 2/5 so for point L (-1, 2) -1 times (2/5) = ? 2 times (2/5) = ? and that would become your new coordenate for L' etc

OpenStudy (anonymous):

dilate means shrink, correct? if so, I (-2,6) times 2/5 = (-0.75, 2.4)

OpenStudy (anonymous):

so since the answer is (-0.75, 2.4), it would round up to (-0.8, 2.4)

jigglypuff314 (jigglypuff314):

no -2 times 2/5 not -0.75 it's -0.8

jigglypuff314 (jigglypuff314):

you shouldn't need to round

OpenStudy (anonymous):

okay